OpenStudy (anonymous):

Solve: m minus three and three-fifths equals four and one-tenth. m equals one and one-half m equals one and seven-tenths m equals seven and one-half m equals seven and seven-tenths

OpenStudy (jim766):

\[m - 3\frac{ 3 }{ 5 }= 4\frac{ 1 }{ 10 }\] is this what it looks like?

OpenStudy (jim766):

can you change these to improper fractions? I hate fractions, I want to get rid of them!

OpenStudy (anonymous):

\[(5*3)+3=\frac{ 18 }{ 5 }\]

OpenStudy (anonymous):

\[(4*10)+1=\frac{ 41 }{ 10}\]

OpenStudy (anonymous):

add 18/5 to 41/10 to get m

OpenStudy (anonymous):

\[\frac{ 41 }{ 10}+\frac{ 15 }{ 8 }=\]

OpenStudy (anonymous):

you can get your answer now @DailynMarie